OpenStudy (anonymous):

2\sqrt{5}+4\sqrt{5}-3\sqrt{5}

jimthompson5910 (jim_thompson5910):

\[\large 2\sqrt{5}+4\sqrt{5}-3\sqrt{5}\] \[\large (2+4-3)\sqrt{5}\] \[\large 3\sqrt{5}\] So \[\large 2\sqrt{5}+4\sqrt{5}-3\sqrt{5}=3\sqrt{5}\]